# Control strategies for magnetized plasma: a polar coordinates framework

### v1

@@
Control strategies for magnetized plasma: a polar coordinates framework
F. Ferrarese Department of Mathematics and Computer Science & Center for Modeling, Computing and Statistics (CMCS), University of Ferrara, via Machiavelli 30, 44121 Ferrara, ITALY (federica.ferrarese@unife.it).
@@ 1 Introduction
In this work, we study the time evolution of plasma density using a Vlasov–Poisson system, which models the dynamics of charged particles under the combined influence of a self-consistent electric field and an externally applied magnetic field. In particular, we focus on the behavior of negatively charged particles, namely electrons, while assuming the ions to form a stationary neutralizing background. Under these assumptions, the governing equations take the form
@@
where the electrons density is defined by
$\rho(t,{\bf x})=\int_{\mathbb{R}^{d_{v}}}f(t,{\bf x},{\bf v})\,d{\bf v}.$ (1.2)
In this formulation, ${\bf E}(t,{\bf x})$, is the self-consistent electric field obtained from the Poisson equation, while ${\bf B}_{ext}(t,{\bf x})$ is an external magnetic field.
@@
where $\mathcal{B}^{adm}$ is a set of admissible controls, $f^{0}$ the initial datum, and $\mathcal{J}(\cdot)$ is a functional which reads as follows
$\mathcal{J}({\bf B}_{ext};f^{0};f)=$ $\int_{0}^{t_{f}}\left(\mathcal{D}(f,\psi)(t)+\frac{\gamma}{2}\int_{\Omega}|{\bf B}_{ext}(t,{\bf x})|^{2}f(t,{\bf x},{\bf v})d{\bf x}d{\bf v}\right)dt$ (1.4)
@@
$\mathcal{D}(f,\psi)(t)=\int_{\Omega}\psi({\bf x},{\bf v})\left(f(t,{\bf x},{\bf v})-\hat{f}(t,{\bf x},{\bf v})\right)d{\bf x}d{\bf v},$ (1.5)
is a running cost function, $\hat{f}$ is a given target distribution discussed next, $\psi({\bf x},{\bf v})$ is a function of the state variables, $\Omega=\Omega_{x}\times\Omega_{v}$, $t_{f}$ is a final prescribed time, and $\gamma$ a weight penalizing the magnitude of the control given by the external magnetic field ${\bf B}_{ext}$. The scope of the functional in (1.4) is to force the distribution function or its moments towards desired values through the choice of the function $\psi({\bf x},{\bf v})$. Many effort have been made in this direction, studying the problem from a theoretical view point as in [3, 34, 5, 32, 41], and from a numerical perspective as in [20, 21]. The aim of the present paper is to provide an overview of the results introduced in [1, 2], reformulating the various control strategies in a two-dimensional framework using polar coordinates. While Cartesian formulations remain widely used, they often fail to exploit the inherent symmetries of confinement devices. In contrast, polar coordinates provide a more natural description of the plasma geometry, especially in Tokamak-like configurations. This choice not only simplifies the representation of magnetic field lines and flux surfaces but also enhances the efficiency and accuracy of numerical discretization. The idea is to construct a piecewise constant control in space by introducing a computational grid composed of cells. In the approach introduced in [1], the control is directly derived as a piecewise constant function within each cell. This strategy is straightforward to implement and ensures that the control can be incorporated efficiently into the numerical scheme at the cell level. In the second approach introduced in [2], the control is initially computed at the level of individual particles, taking into account the local state of the system. The particle-level information is then interpolated to define an average magnetic field within each cell. This approach allows for a finer resolution of the control, as it captures sub-cell variations in the plasma dynamics before producing a cell-averaged control. Consequently, it can provide a more accurate representation of the magnetic field needed to steer the plasma toward the desired configuration, particularly in regions where strong gradients or localized effects are present. Numerical experiments will show a comparison between the two different approaches, proving the efficiency of both strategies.
@@ 2 Problem setting and numerical methods
In this work, we adopt a simplified framework by considering a two-dimensional phase space and investigating the plasma dynamics within a circular domain. We consider a system of particles defined by their position ${\bf x}\in\Omega_{x}\subset\mathbb{R}^{2}$ and velocity ${\bf v}_{\bf x}\in\Omega_{\textrm{v}}\subset\mathbb{R}^{2}$. To better capture the geometry of the setting, we express the dynamics in polar coordinates, assigning to each particle a radial coordinate and an angular coordinate $(r,\theta)\in\Omega_{r}$, such that
@@
where $B(t,{\bf r})>0$. Here and in the following we will use the simplified notation, denoting ${\bf B}_{ext}(\cdot)$ as ${\bf B}(\cdot)$. The two dimensional Vlasov-Poisson equations in polar coordinates reads as
@@
The first equation in (2.8) is designed to enforce a desired configuration of the distribution function corresponding to a prescribed target $\hat{\psi}_{\ell}$. The second equation minimizes the distance from the average value $m(f,\psi_{\ell})$ both in space as well as in velocity space around the target. In the following Sections, we will present an overview on the different control problems introduced in [1, 2], reformulated in polar coordinates.
PIC scheme in polar coordinates
@@
$f^{N}(t^{n+1},{\bf r},{\bf v})=\frac{1}{N}\sum_{m=1}^{N}\delta({\bf r}-{\bf r}_{m}^{n+1})\delta({\bf v}-{\bf v}_{m}^{n+1}),$ (2.9)
@@
$\begin{split}&r_{m}^{n+1}=r_{m}^{n}+hv_{r_{m}}^{n+1},\\ &\theta_{m}^{n+1}=\theta_{m}^{n}+h\frac{v_{\theta_{m}}^{n+1}}{r_{m}^{n+1}},\\ &v_{r_{m}}^{n+1}=v_{r_{m}}^{n}+h\left[E^{n}_{r_{m}}+v_{\theta_{m}}^{n+1}\left(B_{m}^{n}+\frac{v_{\theta_{m}}^{n}}{r_{m}^{n}}\right)\right],\\ &v_{\theta_{m}}^{n+1}=v_{\theta_{m}}^{n+1}+h\left[E_{\theta_{m}}^{n}-v_{r_{m}}^{n+1}\left(B_{m}^{n}+\frac{v_{\theta_{m}}^{n}}{r_{m}^{n}}\right)\right],\end{split}$ (2.11)
@@ 3.1 Strategy one
$f_{k}(t,{\bf r},{\bf v})=\frac{f(t,{\bf r},{\bf v})}{\rho_{k}(t)},\qquad\rho_{k}(t)=\int_{\Omega_{k}}f(t,{\bf r},{\bf v})\,d{\bf r},d{\bf v},$
@@ 3.2 Strategy two
the vector ${\bf B}=[\hat{B}_{1},\ldots,\hat{B}_{N_{c}}]$ collecting the interpolated controls $\hat{B}_{k}$. Here, $\mathcal{I}(\cdot)$ denotes a piecewise-constant interpolation operator, ${\bf r}_{c}$ is the vector of the positions of the centres of the cells $C_{k}$, and ${\bf B}_{N}=[B_{1},\ldots,B_{N}]$ is the vector of pointwise controls $B_{m}$ as defined in equation (3.17). In this framework, the magnetic field acting on a particle located at position ${\bf r}_{m}$ is approximated by the the constant field value assigned to the cell $C_{k}$ containing the particle. While this approach increases the complexity of the control problem, it provides a more realistic and practically implementable setting.
@@ 4 Numerical experiments: Diocotron instability
$\mathcal{E}_{b}^{n}=\frac{1}{2NN_{b}}\sum_{\mathcal{C}_{j}\in\Omega_{b}}\sum_{m=1}^{N}\left|{\bf v}_{m}^{n}-U_{b}^{n}\right|^{2}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j}),$ (4.3)
where
$\begin{split}U_{b}^{n}&=\frac{1}{NN_{b}\rho_{b}^{n}}\sum_{\mathcal{C}_{j}\in\Omega_{b}}\sum_{m=1}^{N}{\bf v}_{m}^{n}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j}),\\ \rho_{b}^{n}&=\frac{1}{NN_{b}}\sum_{\mathcal{C}_{j}\in\Omega_{b}}\sum_{m=1}^{N}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j}),\end{split}$ (4.4)
and $N_{b}$ denotes the number of cells in the space discretization grid, $\mathcal{C}_{j}\in\Omega_{b}$, with $\Omega_{b}=[5-\Delta r,5]\cup[8-\Delta r,8]$. This choice corresponds to a region near the boundaries with width equal to one cell size $\Delta r$.
@@ 4.3 Test with control
Both strategies effectively control the system dynamics, as confirmed by Figure 7, which depicts the evolution of the thermal energy at the boundaries over time. The embedded plot provides a zoomed-in view of the short-time behavior, highlighting the initial decay of the energy. Strategy one results in a slightly greater reduction of boundary thermal energy compared to strategy two.
Figure 7: Diocotron instability: controlled dynamics. Value of the thermal energy in time computed as in (4.3) for both the case where $B(t,{\bf r})$ is chosen as in (3.11) (strategy one), and as in (3.20) (strategy two). The embedded plot provides a zoomed-in view of the short-time behavior.

### v2

@@
Abstract
In this work, we provide an overview of various control strategies aimed at steering plasma toward desired configurations using an external magnetic field. From a modeling perspective, we focus on the Vlasov equation in a two-dimensional bounded domain, accounting for both a self-induced electric field and a strong external magnetic field. The results are presented in a polar coordinate framework, which is particularly well-suited for simulating toroidal devices such as Tokamaks and Stellarators. A key feature of the proposed control strategies is their feedback mechanism, which is based on an instantaneous prediction of the discretized system. Finally, different numerical experiments in the two-dimensional polar coordinate setting demonstrate the effectiveness of the approaches.
@@ 1 Introduction
In this work, we study the time evolution of plasma density using a Vlasov–Poisson system, which models the dynamics of charged particles under the combined influence of a self-consistent electric field and an externally applied magnetic field. In particular, we focus on the behavior of negatively charged particles, namely electrons, while assuming the ions to form a stationary neutral background. Under these assumptions, the equations that govern the evolution of the electron density $f(t,{\bf x},{\bf v})$ take the form
@@
with $({\bf x},{\bf v})\in\mathbb{R}^{2d}$ being the position and velocities, and where the electrons density is defined by
$\rho(t,{\bf x})=\int_{\mathbb{R}^{d_{v}}}f(t,{\bf x},{\bf v})\,d{\bf v},$ (1.2)
while $\rho_{i}$ represents the ion density. In this formulation, ${\bf E}(t,{\bf x})$, is the self-consistent electric field obtained from the Poisson equation as the gradient of the electric potential $\phi(t,{\bf x})$, while ${\bf B}_{ext}(t,{\bf x})$ is an external magnetic field.
@@
where $\mathcal{B}_{adm}$ is a set of admissible controls, $f^{0}$ the initial datum, and $\mathcal{J}(\cdot)$ is a functional which reads as follows
$\mathcal{J}({\bf B}_{ext};f^{0};f)=$ $\int_{0}^{t_{f}}\left(\mathcal{D}(f,\psi)(t)+\frac{\gamma}{2}\int_{\mathbb{R}^{2d}}|{\bf B}_{ext}(t,{\bf x})|^{2}f(t,{\bf x},{\bf v})d{\bf x}d{\bf v}\right)dt$ (1.4)
@@
$\mathcal{D}(f,\psi)(t)=\int_{\mathbb{R}^{2d}}\psi({\bf x},{\bf v})\left(f(t,{\bf x},{\bf v})-\hat{f}(t,{\bf x},{\bf v})\right)d{\bf x}d{\bf v},$ (1.5)
is a running cost function, $\hat{f}$ is a given target distribution discussed next, $\psi({\bf x},{\bf v})$ is a function of the state variables, $t_{f}$ is a final prescribed time, and $\gamma$ a weight penalizing the magnitude of the control given by the external magnetic field ${\bf B}_{ext}$. The scope of the functional in (1.4) is to force the distribution function or its moments towards desired values through the choice of the function $\psi({\bf x},{\bf v})$. Specifically, we aim at reducing the thermal energy and/or the mass at the boundaries of the domain, where excess energy or particle accumulation can create strong gradients or localized instabilities. Many efforts have been made in this direction, studying the problem from a theoretical viewpoint as in [4, 35, 6, 33, 42], and from a numerical perspective as in [21, 22]. The aim of the present paper is to provide an overview of the results introduced in [2, 3], reformulating the various control strategies in a two-dimensional framework using polar coordinates. While Cartesian formulations remain widely used, they often fail to exploit the inherent symmetries of confinement devices. In contrast, polar coordinates provide a more natural description of the plasma geometry, especially in Tokamak-like configurations. This choice not only simplifies the representation of magnetic field lines and flux surfaces but also enhances the efficiency and accuracy of numerical discretization. The idea is to construct a piecewise constant control in space by introducing a computational grid composed of cells. In the approach introduced in [2], the control is directly derived as a piecewise constant function within each cell. This strategy is straightforward to implement and ensures that the control can be incorporated efficiently into the numerical scheme at the cell level. In the second approach introduced in [3], the control is initially computed at the level of individual particles, taking into account the local state of the system. The particle-level information is then interpolated to define an average magnetic field within each cell. This approach allows for a finer resolution of the control, as it captures sub-cell variations in the plasma dynamics before producing a cell-averaged control. Consequently, it can provide a more accurate representation of the magnetic field needed to steer the plasma toward the desired configuration, particularly in regions where strong gradients or localized effects are present. Numerical experiments will show a comparison between the two different approaches, proving the efficiency of both strategies.
@@ 2 Problem setting and numerical methods
In this work, we adopt a simplified framework by considering a two-dimensional phase space. While positions and velocities were initially considered in $\mathbb{R}^{d}$, we now focus on a two-dimensional phase space and investigate the plasma dynamics within a circular spatial domain. Specifically, particles are characterized by their position ${\bf x}\in\Omega_{x}\subset\mathbb{R}^{2}$ and velocity ${\bf v}\in\Omega_{v}\subset\mathbb{R}^{2}$. To better capture the geometry of the setting, we express the dynamics in polar coordinates, assigning to each particle a radial coordinate and an angular coordinate $(r,\theta)\in\Omega_{r}$, such that
@@
where $B(t,{\bf r})>0$. Here and in the following we will use the simplified notation, denoting ${\bf B}_{ext}(\cdot)$ as ${\bf B}(\cdot)$. The two-dimensional Vlasov-Poisson equations in polar coordinates reads as
@@
The first equation in (2.8) is designed to enforce a desired configuration of the distribution function corresponding to a prescribed target $\hat{\psi}_{\ell}$. The second equation minimizes the distance from the average value $m(f,\psi_{\ell})$ both in space as well as in velocity space around the target. In this simplified framework, the goal is to prevent the mass from reaching the upper and lower boundaries of the domain, aiming instead for a particle configuration concentrated near the center with minimal variance.
In the following Sections, we will present an overview on the different control problems introduced in [2, 3], reformulated in polar coordinates.
2.1 PIC scheme in polar coordinates
@@
$f^{N}(t,{\bf r},{\bf v})=\frac{1}{N}\sum_{m=1}^{N}\delta({\bf r}-{\bf r}_{m}(t))\delta({\bf v}-{\bf v}_{m}(t)),$ (2.9)
@@
$\begin{split}&r_{m}^{n+1}=r_{m}^{n}+hv_{r_{m}}^{n+1},\\ &\theta_{m}^{n+1}=\theta_{m}^{n}+h\frac{v_{\theta_{m}}^{n+1}}{r_{m}^{n+1}},\\ &v_{r_{m}}^{n+1}=v_{r_{m}}^{n}+h\left[E^{n}_{r_{m}}+v_{\theta_{m}}^{n+1}\left(B_{m}^{n}+\frac{v_{\theta_{m}}^{n}}{r_{m}^{n}}\right)\right],\\ &v_{\theta_{m}}^{n+1}=v_{\theta_{m}}^{n}+h\left[E_{\theta_{m}}^{n}-v_{r_{m}}^{n+1}\left(B_{m}^{n}+\frac{v_{\theta_{m}}^{n}}{r_{m}^{n}}\right)\right],\end{split}$ (2.11)
@@ 3.1 Strategy one
$f_{k}(t,{\bf r},{\bf v})=\frac{f(t,{\bf r},{\bf v})}{\rho_{k}(t)},\qquad\rho_{k}(t)=\int_{\Omega_{k}}f(t,{\bf r},{\bf v})\,d{\bf r}\,d{\bf v},$
@@ 3.2 Strategy two
the vector ${\bf B}=[\hat{B}_{1},\ldots,\hat{B}_{N_{c}}]$ collecting the interpolated controls $\hat{B}_{k}$. Here, $\mathcal{I}(\cdot)$ denotes a piecewise-constant interpolation operator, ${\bf r}_{c}$ is the vector of the positions of the centres of the cells $C_{k}$, and ${\bf B}_{N}=[B_{1},\ldots,B_{N}]$ is the vector of pointwise controls $B_{m}$ as defined in equation (3.17). In this framework, the magnetic field acting on a particle located at position ${\bf r}_{m}$ is approximated by the constant field value assigned to the cell $C_{k}$ containing the particle. While this approach increases the complexity of the control problem, it provides a more realistic and practically implementable setting.
@@
Remark 2.
Note that the instantaneous controls derived in (3.9)–(3.17) are of order $\mathcal{O}(h^{2})$. To obtain an effective contribution of the control in the limit $h\to 0$, we assume that the parameters scale as in (3.8)-(3.16) with the time step $h$, so that the leading-order is recovered [1]. Additionally, as discussed in Remark 1 of [2], the same instantaneous control laws can be derived directly from the discretized functional over the interval $[t,t+h]$ without explicitly performing any rescaling. This highlights that the scaling is a mathematical strategy to ensure a non-vanishing control, rather than a physical requirement.
@@ 4 Numerical experiments: Diocotron instability
$\mathcal{E}_{b}^{n}=\frac{1}{2N|\Omega_{b}|}\sum_{\mathcal{C}_{j}\in\Omega_{b}}\sum_{m=1}^{N}\left|{\bf v}_{m}^{n}-U_{b}^{n}\right|^{2}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j}),$ (4.3)
where $\chi(\cdot)$ denotes the indicator function and
$\begin{split}U_{b}^{n}&=\frac{1}{N|\Omega_{b}|\rho_{b}^{n}}\sum_{\mathcal{C}_{j}\in\Omega_{b}}\sum_{m=1}^{N}{\bf v}_{m}^{n}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j}),\\ \rho_{b}^{n}&=\frac{1}{N|\Omega_{b}|}\sum_{\mathcal{C}_{j}\in\Omega_{b}}\sum_{m=1}^{N}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j}),\end{split}$ (4.4)
and $\mathcal{C}_{j}\in\Omega_{b}$, with $\Omega_{b}=[5,5+\Delta r]\cup[8-\Delta r,8]$. This choice corresponds to a region near the boundaries with width equal to one cell size $\Delta r$. We also introduce the quantity
$\mathcal{E}_{el}^{n}=\frac{1}{2N|\Omega_{c}|}\sum_{\mathcal{C}_{j}\in\Omega_{c}}\sum_{m=1}^{N}|{\bf E}_{m}^{n}|^{2}\chi({\bf r}_{m}^{n}\in\mathcal{C}_{j})$ (4.5)
which represents the energy associated with the self-induced electric field measured at the center of the domain, namely $\Omega_{c}=[6.5-\Delta r,6.5+\Delta r]$.
@@ 4.3 Test with control
Both strategies effectively control the system dynamics, as confirmed by Figure 7 (left), which depicts the evolution of the thermal energy at the boundaries over time. The embedded plot provides a zoomed-in view of the short-time behavior, highlighting the initial decay of the energy. Strategy one results in a slightly greater reduction of boundary thermal energy compared to strategy two. The application of the controls induces oscillations in the density, particularly in the center of the domain, where regions of high density alternate with regions of low density. This behavior is primarily due to the piecewise-constant structure of the control, which can confine particles at the interfaces of the fictitious macro-cells $C_{k}$, as the control exhibits sharp variations between adjacent cells, particularly in the central region of the domain (see Figure 8). Nevertheless, no increase is observed in the energy associated with the self-induced electric field measured at the center of the domain, as defined in (4.5). We also observe that the electric energy corresponding to strategy two is slightly lower than that obtained with strategy one. This difference may be attributed to the higher particle concentration at the center of the domain produced by strategy one, which leads to a stronger self-induced electric field and, consequently, to an increase in the associated electric energy. This is illustrated in Figure 7 (right), where the electric energy in the uncontrolled case is also shown for comparison.
Figure 7: Diocotron instability. Time evolution of the thermal energy (left) and electric energy (right), computed according to (4.3)–(4.5), for the case in which $B(t,{\bf r})$ is defined as in (3.11) (strategy one) and as in (3.20) (strategy two). The inset in the left panel provides a zoomed-in view of the short-time behavior. In the right panel, the electric energy obtained in the absence of control is also shown for comparison.
@@
Table 1 reports the computational time, in seconds, required for simulations up to the final time $t_{f}=250$, both with and without control, comparing the two proposed strategies. In both cases, the inclusion of the control leads to an increase in computational cost. This is due to the additional operations required, namely identifying the position of each particle within the cells to compute their mean velocity and position in strategy one, or interpolating the pointwise values of the control in strategy two. Strategy two is slightly more computationally demanding than strategy one, which can be attributed to the higher cost of the interpolation procedure. Overall, strategy one achieves a favourable balance between control effectiveness and computational efficiency. All the numerical experiments have been run on a Desktop Computer equipped with an Intel(R) 18 Core(TM) i7-8700 CPU processor and 32GB RAM.
No control Strategy one Strategy two
Time (seconds) 509 1112 1211
Table 1: Diocotron instability. Computational time in seconds of a simulation up to time $t_{f}=250$, without control and with control strategy one and strategy two.
@@ 5 Conclusion
Acknowledges
This work has been written within the activities of GNCS group of INdAM (Italian National Institute of High Mathematics). The research was partially supported by the Italian Ministry of University and Research (MUR) through the PRIN 2020 project (No. 2020JLWP23) ”Integrated Mathematical Approaches to Socio-Epidemiological Dynamics” and funded with the contribution of the Italian Ministry of University and Research (MUR) pursuant to Decree No. 1236 of 01/08/2023 – Fondo Italiano per la Scienza 2022-2023 - Bando FIS 2.